Supposed that you and your friends are in Landmark. Your mother gave you 

P1,000 to buy something for your birthday. You really want to buy bag for 

yourself as a gift, but they are all very expensive. You are now looking for items 

on sale. Then you found one store that has a large sign that reads: all items 

are currently 75% off. This sounds like great news. The original price of the 

bag is ₽1, 500. Does your money enough to buy that bag? What is the fraction 

of the amount of the bag?


1
Expert's answer
2021-09-28T01:26:19-0400
75%100%=34\dfrac{75\%}{100\%}=\dfrac{3}{4}

134=141-\dfrac{3}{4}=\dfrac{1}{4}

Then


14(1500)=375<1000\dfrac{1}{4}(1500)=375<1000

My money enoughs to buy that bag.

New price is 14\dfrac{1}{4} of original price.